# SIC-Free Rate-Splitting Multiple Access: Constellation-Constrained Optimization and Application to Large-Scale Systems

### v1

@@ Abstract
Rate-Splitting Multiple Access (RSMA) has been recognized as a promising multiple access technique for future wireless communication systems. Recent research demonstrates that RSMA can maintain its superiority without relying on Successive Interference Cancellation (SIC) receivers. In practical systems, SIC-free receivers are more attractive than SIC receivers because of their low complexity and latency. This paper evaluates the theoretical limits of RSMA with and without SIC receivers under finite constellations. We first derive the constellation-constrained rate expressions for RSMA. We then design algorithms based on projected subgradient ascent to optimize the precoders and maximize the weighted sum-rate or max-min fairness (MMF) among users. To apply the proposed optimization algorithms to large-scale systems, one challenge lies in the exponentially increasing computational complexity brought about by the constellation-constrained rate expressions. In light of this, we propose methods to avoid such computational burden. Numerical results show that, under optimized precoders, SIC-free RSMA leads to minor losses in weighted sum-rate and MMF performance in comparison to RSMA with SIC receivers, making it a viable option for future implementations.
@@ I Introduction
Most existing work assumes that a Successive Interference Cancellation (SIC) receiver is essential for RSMA, which is referred to as SIC-based RSMA in this work. However, more recent work in [19] initiates research on RSMA without SIC, i.e. SIC-free RSMA [^1] Previous studies use the terms ”non-SIC” and ”SIC-free” interchangeably, as they convey the same meaning., and focuses on its theoretical metric, receiver implementations, and corresponding precoder designs. The findings in [19] suggest that the exclusion of SIC from RSMA receivers reduces complexity, buffer size, and delay and leads to a minor impact on performance. The RSMA implementations discussed in [19] are further verified under the physical layer procedures of the 5G New Radio (NR) and the more practical channel models generated from ray-tracing analysis to support the first industrial contribution to RSMA in 3GPP [20]. Notably, the precoder designs proposed in [19] adopt a low-complexity approach, making it advantageous for practical systems due to its minimal computational burden. Nonetheless, this approach may not fully capture the theoretical limits of SIC-based and SIC-free RSMA. Meanwhile, although RSMA has been evaluated under finite constellations in many works, such as in [19, 9, 21], optimization of RSMA has rarely been carried out utilizing knowledge of finite constellations. This indicates a need to understand the constellation-constrained theoretical limits of RSMA.
@@
Information theory suggests that Gaussian-distributed signal is capacity-achieving for a wide range of communication scenarios. However, it is difficult to apply Gaussian signals in practice because of mainly two reasons: 1) their unbounded amplitude imposes hardware challenges; and 2) their continuous support leads to difficulty in coding and detection. As a result, practical systems often apply input signals with finite alphabets, leading to modulation schemes such as Pulse Amplitude Modulation (PAM) and Quadrature Amplitude Modulation (QAM). It is therefore interesting to investigate the design and performance of communication systems under finite constellations, as they are generally different from those for Gaussian distributed input signals. For example, [22] considers parallel Gaussian channels with arbitrary input distributions and proposes a mercury water-filling policy for power allocation as an extension to the classical water-filling policy for Gaussian inputs. The precoder design for point-to-point MIMO systems is considered in [23, 24]. For uplink multi-user systems with finite constellations, [25] studied two-user multiple access channels with finite constellations, while linear precoding for MIMO multiple access channels was addressed in [26, 27, 28]. For downlink, [29] investigated the precoder design for MIMO broadcast channels. Finite constellations are also considered in more complicated systems, for example, multi-cell systems [30], cognitive radio [31], energy harvesting [32] and relay networks [33]. Although the above-mentioned works focus mainly on mutual information, designs following different criteria, such as mean squared error and diversity, exist and are surveyed in [34]. Note that the computational complexity of mutual information typically increases exponentially with the number of signal sources and cardinality of constellations in use, causing designs for large-scale systems (e.g. massive MIMO) to be difficult and in need of simplification. For example, [26] evaluates the asymptotic rate expression when the number of interference sources tends to be very large. [30] uses Gaussian distribution to approximate the sum of finite alphabet signals that interfere.
@@
  • We derive the constellation-constrained rate expressions and their approximation for SIC-based and SIC-free RSMA. Note that the RSMA rate expressions derived in [19] are based on the assumption that zero-force precoders are used for private streams. As a result of this assumption, inter-user interference was removed during the derivations and the rate expressions were simplified. The rate expressions derived in this work do not make any assumption on the precoders and are therefore more general than those in [19].
  • We propose precoder optimization methods to maximize the Weighted Sum-Rate (WSR) and minimum rate per user (or Max-Min Fairness (MMF)) performance for SIC-based and SIC-free RSMA under finite constellation constraints. To enable this, we developed a global optimal solution to common stream rate allocation for both WSR maximization and MMF problems. These solutions hold for not only finite-constellation inputs but also any other input signal distribution, including Gaussian inputs.
  • We observe that the computational complexity of the proposed algorithms grows exponentially as the system scales up in terms of the number of antennas, the number of users, and the cardinality of constellations. In light of this, we propose methods to scale up optimization algorithms to handle large-scale systems and avoid prohibitive computational cost.
  • We show through simulations that both SIC-based and SIC-free RSMA outperform SDMA under different evaluation metrics and dimensions of the system. Moreover, SIC-free RSMA often preserves most of the superiority of SIC-based RSMA. The results also show that SIC-based and SIC-free RSMA require very different precoders, despite the fact that they achieve similar performance in the end. This indicates the need for dedicated precoder designs for SIC-free RSMA. Therefore, we extend the conclusion in [19] from low-complexity precoders to fully optimized precoders, from sum-rate objective to WSR and MMF objectives, and from small-scale to large-scale systems.
Organization: The rest of this paper is organized as follows. Section II introduces a system model for RSMA with explanations on SIC and non-SIC receivers. In Section III, derivation of achievable rate expressions of RSMA under finite constellations is provided. In Section IV, precoder optimization algorithms for WSR and MMF optimization are proposed. Section V addresses optimization in large-scale systems. Section VI presents and analyzes numerical results. Finally, Section VII concludes this paper.
@@ II RSMA System Model
Figure 1: Single-layer RSMA with SIC and non-SIC receivers.
The transmitter applies single-layer RSMA [6] as the multiple access scheme, whose key procedures are shown in Figure 1 and are explained as follows. Let $\{W_{1},...,W_{K}\}$ respectively denote the $K$ unicast messages to be transmitted to the $K$ users. The transmitter first processes the messages by splitting each message into common and private parts, e.g. $W_{k}$, $\forall k\in\mathcal{K}$ is split into $W_{\text{c},k}$ and $W_{\text{p},k}$. Next, the $K$ common parts, i.e. $\{W_{\text{c},1},...,W_{\text{c},K}\}$ are combined into one common message, denoted by $W_{\text{c}}$, while the private parts, i.e. $\{W_{\text{p},1},...,W_{\text{p},K}\}$ are treated as individual messages. The resulting $K+1$ messages, $\{W_{\text{c}},W_{\text{p},1},...,W_{\text{p},K}\}$, are then encoded in $K+1$ common and private symbol streams, denoted by $\{s_{\text{c}},s_{1},...,s_{K}\}$. Finally, these symbols are precoded and transmitted. Therefore, we express the transmitted signal as
@@
where $\mathbf{P}=[\mathbf{p}_{\text{c}},\mathbf{p}_{1},...,\mathbf{p}_{K}]$ denotes the precoding matrix for the common and private streams, and $\mathbf{s}=[s_{\text{c}},s_{1},...,s_{K}]^{T}$. We assume that the symbols are drawn independently and uniformly from finite constellations with zero mean and unit average power. Therefore, $\mathsf{E}\{\mathbf{s}\mathbf{s}^{H}\}=\mathbf{I}$.
@@
  • SIC receiver: The receiver first decodes $W_{\text{c}}$ from $y_{k}$ by treating $\{s_{1},...,s_{K}\}$ as noise and obtains $\hat{W}_{\text{c}}$, denoting the resulting estimate of $W_{\text{c}}$. $\hat{W}_{\text{c}}$ is then re-encoded into $\hat{s}_{\text{c}}$, an estimate of the common stream symbol. Next, the receiver decodes $W_{\text{p},k}$ from $y_{k}-\mathbf{h}_{k}^{H}\mathbf{p}_{\text{c}}\hat{s}_{\text{c}}$, i.e. the received signal with the estimated common stream symbol canceled, by treating $\{s_{k^{\prime}}\}$, $\forall k^{\prime}\in\mathcal{K}/k$, as noise and obtains $\hat{W}_{\text{p},k}$, an estimate of the desired private message. Finally, the receiver extracts $\hat{W}_{\text{c},k}$ from $\hat{W}_{\text{c}}$ and combines it with $\hat{W}_{\text{p},k}$ to reconstruct $\hat{W}_{k}$ as the resulting estimate of the original unicast message for user-$k$.
@@
    Hence, $\hat{W}_{\text{c}}=W_{\text{c}}$, i.e. the cancellation of common stream is always perfect.
  • Non-SIC receiver: The receiver decodes $W_{\text{c}}$ from $y_{k}$ by treating $\{s_{1},...,s_{K}\}$ as noise, and also decodes $W_{\text{p},k}$ by treating $s_{\text{c}}$ and $s_{k}^{\prime}$, $\forall k^{\prime}\in\mathcal{K}/k$, as noise. After this, the receiver extracts $\hat{W}_{\text{c},k}$ from $\hat{W}_{\text{c}}$ and combines it with $\hat{W}_{\text{p},k}$ to reconstruct $\hat{W}_{k}$. The main difference from the SIC implementation is that $\hat{s}_{\text{c}}$ remains in the received signal when the receiver decodes $W_{\text{p},k}$ due to the absence of the cancellation step.
@@ III-A Prerequisites on Constellation-Constrained Rate
where $n\sim\mathcal{CN}(0,\sigma^{2})$, $\mathbf{h}^{H}\mathbf{P}$ is fixed and known, and $\mathbf{s}$ is an observation of a uniformly distributed random variable on finite support given by $\mathcal{X}$. Let $\mathbf{s}_{m}$ denote the $m$-th element in $\mathcal{X}$ in a certain order. The mutual information in bits between $\mathbf{s}$ and $y$ is given by
@@
$\begin{split}&H(\mathbf{s}|y,\mathbf{h}^{H}\mathbf{P})\\ &=\frac{1}{\ln 2}+\frac{1}{|\mathcal{X}|}\sum_{m=1}^{|\mathcal{X}|}\mathsf{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}|}\exp\Big{(}-\frac{|\mathbf{h}^{H}\mathbf{P}\mathbf{d}_{m,l}(\mathcal{X})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}},\end{split}$ (5)
@@ III-B Achievable Rate of RSMA
$R_{\text{c},k}=$ $\log_{2}|\mathcal{X}_{\text{c}}|-\frac{1}{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\sum_{m=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\mathsf{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\mathbf{h}^{H}\mathbf{p}_{\text{c}}\mathcal{X}_{\text{c}}+\sum_{a\in\mathcal{K}}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}}$
$+\frac{1}{|\mathcal{X}_{\mathcal{K}}|}\sum_{m=1}^{|\mathcal{X}_{\mathcal{K}}|}\mathsf{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\mathcal{K}}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\sum_{a\in\mathcal{K}}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}}$ (7)
$R_{\text{p},k}^{\text{SIC}}=$ $\log_{2}|\mathcal{X}_{k}|-\frac{1}{|\mathcal{X}_{\mathcal{K}}|}\sum_{m=1}^{|\mathcal{X}_{\mathcal{K}}|}\mathsf{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\mathcal{K}}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\sum_{a\in\mathcal{K}}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}}$
$+\frac{1}{|\mathcal{X}_{\mathcal{K}/k}|}\sum_{m=1}^{|\mathcal{X}_{\mathcal{K}/k}|}\mathsf{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\mathcal{K}/k}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\sum_{a\in\mathcal{K}/k}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}}$ (8)
$R_{\text{p},k}^{\text{non-SIC}}=$ $\log_{2}|\mathcal{X}_{k}|-\frac{1}{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\sum_{m=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\mathsf{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\mathbf{h}^{H}\mathbf{p}_{\text{c}}\mathcal{X}_{\text{c}}+\sum_{a\in\mathcal{K}}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}}$
$+\frac{1}{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}/k}|}\sum_{m=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}/k}|}\mathsf{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}/k}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\mathbf{h}^{H}\mathbf{p}_{\text{c}}\mathcal{X}_{\text{c}}+\sum_{a\in\mathcal{K}/k}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}}$ (9)
@@
where the expansion is made from the chain rule of mutual information [36]. To ensure that the common stream can be decoded by all the users, the common stream rate depends on the worst-cast user, i.e.
@@
$\begin{split}&R_{\text{p},k}^{\text{non-SIC}}\\ &=I\Big{(}s_{k};y_{k}\Big{|}\mathbf{h}_{k}^{H}\mathbf{P}\Big{)}\\ &=H(s_{\text{c}},\{s_{k}|k\in\mathcal{K}\})-H\Big{(}s_{\text{c}},\{s_{k}|k\in\mathcal{K}\}\Big{|}y_{k},\mathbf{h}_{k}^{H}\mathbf{P}\Big{)}\\ &\;\;\;\;-H(s_{\text{c}},\{s_{k^{\prime}}|k^{\prime}\in\mathcal{K}/k\})\\ &\;\;\;\;+H\Big{(}s_{\text{c}},\{s_{k^{\prime}}|k^{\prime}\in\mathcal{K}/k\}\Big{|}y_{k}-\mathbf{h}_{k}^{H}\mathbf{p}_{k}s_{k},\mathbf{h}_{k}^{H}\mathbf{P}\Big{)}.\end{split}$ (13)
Note that the difference between (12) and (13) lies in the availability of knowledge on $s_{\text{c}}$. The full expressions of $R_{\text{c},k}$, $R_{\text{p},k}^{\text{SIC}}$ and $R_{\text{p},k}^{\text{non-SIC}}$ are given by (III-B)-(III-B) respectively, where $\mathcal{X}_{\text{c}}$ is the constellation set used for the common stream, $\mathcal{X}_{\text{p},k}$ is the constellation set used for the private stream intended for the $k$-th user, and $\mathcal{X}_{\mathcal{M}}\triangleq\prod_{k\in\mathcal{M}}\mathcal{X}_{\text{p},k}$ for certain set $\mathcal{M}$. The approximations of $R_{\text{c},k}$, $R_{\text{p},k}^{\text{SIC}}$ and $R_{\text{p},k}^{\text{non-SIC}}$ are given by (III-B)-(III-B) respectively.

### v2

@@ Abstract
Rate-Splitting Multiple Access (RSMA) has been recognized as a promising multiple access technique for future wireless communication systems. Recent research demonstrates that RSMA can maintain its superiority without relying on Successive Interference Cancellation (SIC) receivers. In practical systems, SIC-free receivers are more attractive than SIC receivers because of their low complexity and latency. This paper evaluates the theoretical limits of RSMA with and without SIC receivers under finite constellations. We first derive the constellation-constrained rate expressions for RSMA. We then design algorithms based on projected subgradient ascent to optimize the precoders and maximize the weighted sum-rate or max-min fairness among users. To apply the proposed optimization algorithms to large-scale systems, one challenge lies in the exponentially increasing computational complexity brought about by the constellation-constrained rate expressions. In light of this, we propose methods to avoid such computational burden. Numerical results show that, under optimized precoders, SIC-free RSMA leads to minor losses in both weighted sum-rate and max-min fairness in comparison to RSMA with SIC receivers, making it a viable option for future implementations.
@@ I Introduction
Most existing work assumes that a Successive Interference Cancellation (SIC) receiver is essential for RSMA, which is referred to as SIC-based RSMA in this work. However, more recent work in [19] initiates research on RSMA without SIC, i.e., SIC-free RSMA111[1] and [19] use the terms ”SIC-free” and ”non-SIC” interchangeably, as they convey the same meaning. In this work, we adopt the term ”SIC-free” for consistency, and refer ”RSMA with SIC receivers” as ”SIC-based RSMA” for clarity., and focuses on its theoretical metric, receiver implementations, and the corresponding precoder designs. The findings in [19] suggest that the exclusion of SIC from RSMA receivers reduces complexity, buffer size, and delay and leads to a minor impact on performance. The RSMA implementations discussed in [19] are further verified under the physical layer procedures of the 5G New Radio (NR) and the more practical channel models generated from ray-tracing analysis to support the first industrial contribution to RSMA in 3GPP [20]. Notably, the precoder designs proposed in [19] adopt a low-complexity approach, making it advantageous for practical systems due to its minimal computational burden. Nonetheless, this approach may not fully capture the theoretical limits of SIC-based and SIC-free RSMA. Meanwhile, although RSMA has been evaluated under finite constellations in many works, such as in [19, 9, 21], optimization of RSMA has rarely been carried out utilizing knowledge of finite constellations. This indicates a need to understand the constellation-constrained theoretical limits of RSMA.
@@
Information theory suggests that Gaussian-distributed input signal is capacity-achieving for a wide range of communication scenarios. However, it is difficult to apply Gaussian signals in practice because of mainly two reasons: 1) their unbounded amplitude imposes hardware challenges; and 2) their continuous support leads to difficulty in coding and detection. As a result, practical systems often apply input signals with finite alphabets, leading to modulation schemes such as Pulse Amplitude Modulation (PAM) and Quadrature Amplitude Modulation (QAM). It is therefore interesting to investigate the design and performance of communication systems under finite constellations, as they are generally different from those for Gaussian distributed input signals. For example, [22] considers parallel Gaussian channels with arbitrary input distributions and proposes a mercury water-filling policy for power allocation as an extension to the classical water-filling policy for Gaussian inputs. The precoder design for point-to-point MIMO systems is considered in [23, 24]. For uplink multi-user systems with finite constellations, [25] studied two-user multiple access channels with finite constellations, while linear precoding for MIMO multiple access channels was addressed in [26, 27, 28]. For downlink, [29] investigated the precoder design for MIMO broadcast channels. Finite constellations are also considered in more complicated systems, for example, multi-cell systems [30], cognitive radio [31], energy harvesting [32] and relay networks [33]. Although the above-mentioned works focus mainly on mutual information, designs following different criteria, such as mean squared error and diversity, exist and are surveyed in [34]. Note that the computational complexity of mutual information typically increases exponentially with the number of signal sources and the cardinality of constellations in use, causing designs for large-scale systems (e.g., massive MIMO) to be difficult and in need of simplification. For example, [26] evaluates the asymptotic rate expression when the number of interference sources tends to be very large. [30] uses Gaussian distribution to approximate the sum of finite alphabet signals that interfere.
@@
  • We derive the constellation-constrained rate expressions and their approximation for SIC-based and SIC-free RSMA. Note that the RSMA rate expressions derived in [19] are based on the assumption that zero-forcing precoders are used for private streams. As a result of this assumption, inter-user interference was removed during the derivations and the rate expressions were simplified. The rate expressions derived in this work do not make any assumption on the precoders and are therefore more general than those in [19].
  • We propose precoder optimization methods to maximize the Weighted Sum-Rate (WSR) and Max-Min Fairness (MMF) (or minimum rate per user) performance for SIC-based and SIC-free RSMA under finite constellation constraints. To enable this, we developed global optimal solutions to common stream rate allocation for both WSR maximization and MMF problems. These solutions hold for not only finite-constellation inputs but also any other input signal distribution, including Gaussian inputs.
  • We observe that the computational complexity of the proposed algorithms grows exponentially as the system scales up in terms of the number of users. In light of this, we propose methods to scale up optimization algorithms to handle large-scale systems and avoid prohibitive computational cost.
  • We show through simulations that both SIC-based and SIC-free RSMA outperform SDMA under different evaluation metrics and dimensions of the system. Moreover, SIC-free RSMA often preserves most of the superiority of SIC-based RSMA. The results also show that SIC-based and SIC-free RSMA require very different precoders, despite the fact that they achieve similar performance in the end. This indicates the need for dedicated precoder designs for SIC-free RSMA. Therefore, we extend the conclusion in [19] from low-complexity precoders to fully optimized precoders, from the sum-rate objective to the WSR and MMF objectives, and from small-scale to large-scale systems.
Organization: The rest of this paper is organized as follows. Section II introduces a system model for RSMA with explanations of SIC and SIC-free receivers. In Section III, derivation of achievable rate expressions for RSMA under finite constellations is provided. In Section IV, precoder optimization algorithms for WSR and MMF optimization are proposed. Section V addresses optimization in large-scale systems. Section VI presents and analyzes numerical results. Finally, Section VII concludes this paper.
@@ II RSMA System Model
Figure 1: Single-layer RSMA with SIC and SIC-free receivers.
The transmitter applies single-layer RSMA [6] as the multiple access scheme, whose key procedures are shown in Figure 1 and are explained as follows. Let $\{W_{1},...,W_{K}\}$ respectively denote the $K$ unicast messages to be transmitted to the $K$ users. The transmitter first processes the messages by splitting each message into common and private parts, e.g., $W_{k}$, $\forall k\in\mathcal{K}$, is split into $W_{\text{c},k}$ and $W_{\text{p},k}$. Next, the $K$ common parts, i.e., $\{W_{\text{c},1},...,W_{\text{c},K}\}$, are combined into one common message, denoted by $W_{\text{c}}$, while the private parts, i.e., $\{W_{\text{p},1},...,W_{\text{p},K}\}$, are treated as individual messages. The resulting $K+1$ messages, $\{W_{\text{c}},W_{\text{p},1},...,W_{\text{p},K}\}$, are then encoded in $K+1$ common and private symbol streams, denoted by $\{s_{\text{c}},s_{1},...,s_{K}\}$. Finally, these symbols are precoded and transmitted. Therefore, we express the transmitted signal as
@@
where $\mathbf{P}=[\mathbf{p}_{\text{c}},\mathbf{p}_{1},...,\mathbf{p}_{K}]$ denotes the precoding matrix for the common and private streams, and $\mathbf{s}=[s_{\text{c}},s_{1},...,s_{K}]^{T}$. We assume that the symbols are drawn independently and uniformly from finite constellations with zero mean and unit average power. Therefore, $\mathbb{E}\{\mathbf{s}\mathbf{s}^{H}\}=\mathbf{I}^{K+1}$.
@@
  • SIC receiver: The receiver first decodes $W_{\text{c}}$ from $y_{k}$ by treating $\{s_{1},...,s_{K}\}$ as noise and obtains $\hat{W}_{\text{c}}$, denoting the resulting estimate of $W_{\text{c}}$. $\hat{W}_{\text{c}}$ is then re-encoded into $\hat{s}_{\text{c}}$, an estimate of the common stream symbol. Next, the receiver decodes $W_{\text{p},k}$ from $y_{k}-\mathbf{h}_{k}^{H}\mathbf{p}_{\text{c}}\hat{s}_{\text{c}}$, i.e., the received signal with the estimated common stream symbol canceled, by treating $s_{k^{\prime}}$, $\forall k^{\prime}\in\mathcal{K}/k$, as noise and obtains $\hat{W}_{\text{p},k}$, an estimate of the desired private message. Finally, the receiver extracts $\hat{W}_{\text{c},k}$ from $\hat{W}_{\text{c}}$ and combines it with $\hat{W}_{\text{p},k}$ to reconstruct $\hat{W}_{k}$ as the resulting estimate of the original unicast message for user-$k$.
@@
    Hence, $\hat{W}_{\text{c}}=W_{\text{c}}$, i.e., the cancellation of common stream is always perfect.
  • SIC-free receiver: The receiver decodes $W_{\text{c}}$ from $y_{k}$ by treating $\{s_{1},...,s_{K}\}$ as noise, and also decodes $W_{\text{p},k}$ by treating $s_{\text{c}}$ and $s_{k}^{\prime}$, $\forall k^{\prime}\in\mathcal{K}/k$, as noise. After this, the receiver extracts $\hat{W}_{\text{c},k}$ from $\hat{W}_{\text{c}}$ and combines it with $\hat{W}_{\text{p},k}$ to reconstruct $\hat{W}_{k}$. The main difference from the SIC implementation is that $\hat{s}_{\text{c}}$ remains in the received signal when the receiver decodes $W_{\text{p},k}$ due to the absence of the cancellation step.
@@ III-A Prerequisites on Constellation-Constrained Rate
where $n\sim\mathcal{CN}(0,\sigma^{2})$, $\mathbf{h}^{H}\mathbf{P}$ is fixed and known, and $\mathbf{s}$ is an observation of a uniformly distributed random variable with finite support given by $\mathcal{X}$. Let $\mathbf{s}_{m}$ denote the $m$-th element in $\mathcal{X}$ in a certain order. The mutual information in bits between $\mathbf{s}$ and $y$ is given by
@@
$\begin{split}&H(\mathbf{s}|y,\mathbf{h}^{H}\mathbf{P})\\ &=\frac{1}{\ln 2}+\frac{1}{|\mathcal{X}|}\sum_{m=1}^{|\mathcal{X}|}\mathbb{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}|}\exp\Big{(}-\frac{|\mathbf{h}^{H}\mathbf{P}\mathbf{d}_{m,l}(\mathcal{X})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}},\end{split}$ (5)
@@ III-B Achievable Rate of RSMA
$R_{\text{c},k}=$ $\log_{2}|\mathcal{X}_{\text{c}}|-\frac{1}{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\sum_{m=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\mathbb{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\mathbf{h}^{H}\mathbf{p}_{\text{c}}\mathcal{X}_{\text{c}}+\sum_{a\in\mathcal{K}}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}}$
$+\frac{1}{|\mathcal{X}_{\mathcal{K}}|}\sum_{m=1}^{|\mathcal{X}_{\mathcal{K}}|}\mathbb{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\mathcal{K}}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\sum_{a\in\mathcal{K}}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}}$ (7)
$R_{\text{p},k}^{\text{SIC}}=$ $\log_{2}|\mathcal{X}_{k}|-\frac{1}{|\mathcal{X}_{\mathcal{K}}|}\sum_{m=1}^{|\mathcal{X}_{\mathcal{K}}|}\mathbb{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\mathcal{K}}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\sum_{a\in\mathcal{K}}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}}$
$+\frac{1}{|\mathcal{X}_{\mathcal{K}/k}|}\sum_{m=1}^{|\mathcal{X}_{\mathcal{K}/k}|}\mathbb{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\mathcal{K}/k}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\sum_{a\in\mathcal{K}/k}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}}$ (8)
$R_{\text{p},k}^{\text{SIC-free}}=$ $\log_{2}|\mathcal{X}_{k}|-\frac{1}{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\sum_{m=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\mathbb{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\mathbf{h}^{H}\mathbf{p}_{\text{c}}\mathcal{X}_{\text{c}}+\sum_{a\in\mathcal{K}}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}}$
$+\frac{1}{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}/k}|}\sum_{m=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}/k}|}\mathbb{E}_{n}\Big{\{}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}/k}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\mathbf{h}^{H}\mathbf{p}_{\text{c}}\mathcal{X}_{\text{c}}+\sum_{a\in\mathcal{K}/k}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})+n|^{2}}{\sigma^{2}}\Big{)}\Big{\}}$ (9)
@@
where the expansion is made from the chain rule of mutual information [36]. To ensure that the common stream can be decoded by all the users, the common stream rate depends on the worst-case user, i.e.,
@@
$\begin{split}&R_{\text{p},k}^{\text{SIC-free}}\\ &=I\Big{(}s_{k};y_{k}\Big{|}\mathbf{h}_{k}^{H}\mathbf{P}\Big{)}\\ &=H(s_{\text{c}},\{s_{k}|k\in\mathcal{K}\})-H\Big{(}s_{\text{c}},\{s_{k}|k\in\mathcal{K}\}\Big{|}y_{k},\mathbf{h}_{k}^{H}\mathbf{P}\Big{)}\\ &\;\;\;\;-H(s_{\text{c}},\{s_{k^{\prime}}|k^{\prime}\in\mathcal{K}/k\})\\ &\;\;\;\;+H\Big{(}s_{\text{c}},\{s_{k^{\prime}}|k^{\prime}\in\mathcal{K}/k\}\Big{|}y_{k}-\mathbf{h}_{k}^{H}\mathbf{p}_{k}s_{k},\mathbf{h}_{k}^{H}\mathbf{P}\Big{)}.\end{split}$ (13)
Note that the difference between (12) and (13) lies in the availability of knowledge on $s_{\text{c}}$. The full expressions of $R_{\text{c},k}$, $R_{\text{p},k}^{\text{SIC}}$ and $R_{\text{p},k}^{\text{SIC-free}}$ are given by (III-B)-(III-B) respectively, where $\mathcal{X}_{\text{c}}$ is the constellation set used for the common stream, $\mathcal{X}_{\text{p},k}$ is the constellation set used for the private stream intended for the $k$-th user, and $\mathcal{X}_{\mathcal{M}}\triangleq\prod_{k\in\mathcal{M}}\mathcal{X}_{\text{p},k}$ for certain set $\mathcal{M}$. The approximations of $R_{\text{c},k}$, $R_{\text{p},k}^{\text{SIC}}$ and $R_{\text{p},k}^{\text{SIC-free}}$ are given by (III-B)-(III-B) respectively.